I need to prepare a 5M solution of NaCl. How would I do that? Then I'd need to use that solution to prepare some other 5L volume reagent with a final concentration of 1M NaCl. How would I make that?

Tell me about the role of magnesium in dna replication.
Answer #1

For preparing 5M NaCl solution, Since 1M is the molecular weight mass in grams one litre of solvent. Hence the molecular weight of NaCl is 58.44 g/mol. For 5M NaCl solution 58.44 * 5 = 292.2 g dissolved in 1 litre of distilled water.

For Preparing 5 litre solution with 1M Final NaCl concentration

M1V1 = M2V2, 5M * V1 = 5L * 1M (M1 = initial conc, M2 final conc, V1 required volume, V2 final volume to be made)

= 1 Litre of 5M NaCl solution to be mixed for making your required solution.

2. DNA replication is a cofactor for DNA polymerase, it facilitates the formation of nucleotide monophosphates from tri phosphates. Hence the insertion of correct nucleotide and further its conversion to monophosphate.